Math

You put $1 in the bank for one year. You will

get 100% annual interest. Find the amount of the 1 year annually, semiannually, quarterly, monthly, and weekly.
Answer #1

Annually: 1 * (1 + 1/1)1 = $2 

Semiannually: 1 * (1 + 1/2)2 * 1 = $2.25 

Quarterly: 1 * (1 + 1/4)4 * 1 = $2.441... 

Monthly: 1 * (1 + 1/12)12 * 1 = $2.613... 

Weekly: 1 * (1 + 1/52)52 * 1 = $2.69257

answered by: Tina Gao